OpenStudy (anonymous):

-3(a-2)+1=28 , howw yu got the answerr

OpenStudy (anonymous):

-3a+6+1=28.....-3a=21....a=-7

OpenStudy (anonymous):

-3(a-2)+1=28 (distribute the -3 to a and -2) -3a + 6 + 1 = 28 (add 6 and 1) -3a + 7 = 28 (subtract 7 from each side) -7 -7 -3a = 21 (divide -3 from both sides) /-3 /-3 a = -7

OpenStudy (anonymous):

-3(a-2)+1=28 =>-3a+6+1=28 =>-3a +7 =28 =>-3a=21 =>a=21/-3 =>a=-7